# RAG+: Enhancing Retrieval-Augmented Generation with Application-Aware Reasoning

### v3

@@ Abstract
The integration of external knowledge through Retrieval-Augmented Generation (RAG) has become foundational in enhancing large language models (LLMs) for knowledge-intensive tasks. However, existing RAG paradigms often overlook the cognitive step of applying knowledge, leaving a gap between retrieved facts and task-specific reasoning. In this work, we introduce RAG+, a principled and modular extension that explicitly incorporates application-aware reasoning into the RAG pipeline. RAG+ constructs a dual corpus consisting of knowledge and aligned application examples, created either manually or automatically, and retrieves both jointly during inference. This design enables LLMs not only to access relevant information but also to apply it within structured, goal-oriented reasoning processes. Experiments across mathematical, legal, and medical domains, conducted on multiple models, demonstrate that RAG+ consistently outperforms standard RAG variants, achieving average improvements of 3–5%, and peak gains up to 7.5% in complex scenarios. By bridging retrieval with actionable application, RAG+ advances a more cognitively grounded framework for knowledge integration, representing a step toward more interpretable and capable LLMs.
footnotetext: * Work done during an internship at Huawei Technologies Ltd.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a broad range of natural language processing tasks (Rong et al., 2025; Tang et al., 2025). To further enhance their capabilities, Retrieval-Augmented Generation (RAG) has become a widely adopted framework. By equipping LLMs with access to external knowledge sources, RAG enables the dynamic retrieval of up-to-date information at inference time, significantly improving performance in knowledge-intensive scenarios (Li et al., 2025; Mostafa et al., 2025).
@@
While effective in factual recall and open-domain question answering, RAG frequently underperforms on domain-specific reasoning tasks (Lin et al., 2025; Hayashi et al., 2025; Ammann et al., 2025), where solving complex problems requires not only relevant information but also reasoning about how to use it to arrive at a solution, as shown in Figure 1.
RA-DIT Lin et al. (2023) fine-tunes the retriever and generator in a dual-instruction manner, aligning retrieval more closely with what the model needs to generate accurate responses. Other recent RAG extensions tackle this issue by decomposing reasoning tasks into smaller steps and retrieving relevant knowledge for each substep (Singh et al., 2025; Xiong et al., 2025; Zihao et al., 2024). However, they provide limited guidance on applying retrieved knowledge, which hampers performance in procedural reasoning tasks that require understanding both the process and the underlying concepts.
@@
Motivated by these insights, we propose RAG+, a simple yet effective extension to the RAG framework that enhances reasoning by bridging retrieval and generation with an application-aware stage. Instead of retrieving only relevant knowledge, RAG+ additionally retrieves examples that demonstrate how the knowledge is applied in practice—such as structured reasoning chains and stepwise solutions—to ground the model’s output in task-relevant usage and to improve reasoning accuracy.
RAG+ builds a dual corpus: one stores domain knowledge, and the other containing automatically generated application instances aligned to each fact. At inference time, the system retrieves relevant knowledge based on the input query and then fetches aligned application examples to provide practical context. This design encourages the model not only to recall facts but also to produce outputs that follow grounded reasoning patterns based on prior usage. RAG+ is modular and retrieval-agnostic; it can be integrated into any existing RAG pipeline without changing the model architecture or requiring additional fine-tuning.
@@
  • We identify a key limitation of existing RAG systems: the lack of an application-aware step that impeding reasoning in complex tasks.
@@ 2 Related Work
For retrieval quality, Rewrite-Retrieve-Read (Ma et al., 2023) rewrites queries to better match document style. $R^{2}AG$ (Ye et al., 2024) uses a retrieval-aware encoder to highlight key signals, while Query2Doc (Wang et al., 2023a) expands queries into pseudo-documents to clarify intent. On the corpus side, GraphRAG (Edge et al., 2024) integrates knowledge graphs to support entity-centric reasoning. To enhance retrieval precision, Reranking (Wang et al., 2024a) and filtering techniques (Wang et al., 2023b; Pickett et al., 2025) further refine retrieved content.
@@ 3.1 Construction Stage
Application Matching: In domains where real-world cases naturally exemplify the use of specific knowledge, each knowledge item is paired with one or more application instances drawn from authentic scenarios, serving as grounded demonstrations that contextualize and concretize the corresponding knowledge. To establish these pairings, we first perform category alignment by assigning both problems and knowledge items to broad categories using powerful LLMs with temperature sampling and self-consistency voting, followed by manual refinement to ensure accuracy. Next, within each category, relevance selection is conducted by prompting the model to identify the most pertinent knowledge entries for each problem.
@@ 3.2 Inference Stage
During inference, given a test query, RAG+ first retrieves relevant knowledge items from the knowledge corpus using any retrieval method (e.g., dense retrieval, reranking), as shown in Figure 2(b). For each retrieved knowledge item $k$, its corresponding application example $a$, pre-aligned during the construction stage, is retrieved from the application corpus. The pair $(k,a)$ is then incorporated into a predefined prompt template that guides the model with both factual information and procedural cues. This prompt is subsequently fed into the language model for final answer generation.
@@
Methods GLM4-9B Qwen2.5-7B DS-Qwen-7B Qwen2.5-14B DS-Qwen-32B Qwen2.5-72B LLaMA3.3-70B
@@ 4.3 Models
Methods LLaMA3.1-8B* DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen3-32B Qwen2.5-72B LLaMA3.3-70B
@@ 5 Results
Methods LLaMA3.1-8B Qwen2.5-7B DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen2.5-72B LLaMA3.3-70B
@@ 5.4 Impact of Reranking Model
Methods Qwen2.5-7B Qwen2.5-14B
@@
Methods Qwen2.5-7B Qwen2.5-14B
@@ 5.5 Effect of Application Examples Only
We conducted experiments to assess the contribution of application examples without their associated knowledge on the legal domain. Results are shown in Table 6. Incorporating only application examples leads to performance improvements over the base models in most cases. This indicates that application-level guidance alone provides a meaningful benefit. However, this setting still underperforms compared to the full configuration where both knowledge and application examples are included, referred to as RAG+ and AFRAG+.
@@
Methods Qwen2.5-7B Qwen2.5-14B DS-Qwen-32B
@@ 6 Discussion
To enable application-aware retrieval, we propose a lightweight method for constructing application-specific corpora from a knowledge base, typically generating one to two examples per item. The corpus grows linearly with the number of items and introduces no retrieval overhead, as each item is directly paired with an application example. In experiments, the initial corpus sizes were 223 KB (math), 528 KB (legal), and 99,382 KB (medicine). Generating the legal corpus with Qwen2.5-72B on eight 64 GB NPUs took around six hours, which is acceptable given this scale. Final sizes reached 612 KB, 868 KB, and 105,558 KB, respectively, closely matching the sizes of the underlying knowledge corpora. The method is scalable and supports efficient incremental updates.
@@ Appendix A Experimental Details
For models generating long-form outputs (i.e., QwQ-32B, DS-Qwen-7B, and DS-Qwen-32B), we set the decoding temperature to 1 and top_p set to 1 to encourage output diversity. For all other models, we apply a deterministic decoding strategy with a temperature of 0 and top_p of 1. Each experiment is conducted three times, and we report the average performance across these runs.
@@ A.3.1 Math QA in Mathematic Domain
We first collected a range of publicly available online resources, supplemented with original questions and corresponding answers. The dataset includes both standard QA items and multiple-choice questions with varying numbers of answer options. Only the correct answers were retained. For compound questions that inquire about multiple values (e.g., "What is the value of $y$ when $x=0$ or $x=1$?"), we decomposed them into separate single-answer questions, such as "What is the value of $y$ when $x=0$" and "What is the value of $y$ when $x=1$?".
Next, we transformed all QA pairs into a multiple-choice format by prompting GPT to generate three additional plausible but incorrect answer options. This step expands the test set while simultaneously constraining the model’s generative space.
@@ B.1 Complete Results
Methods LLaMA3.1-8B* DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen3-32B Qwen2.5-72B LLaMA3.3-70B
@@
Methods LLaMA3.1-8B Qwen2.5-7B DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen2.5-72B LLaMA3.3-70B
@@
Methods GLM4-9B Qwen2.5-7B DS-Qwen-7B Qwen2.5-14B DS-Qwen-32B Qwen2.5-72B LLaMA3.3-70B
@@
Methods Qwen2.5-14B Qwen3-32B DS-Qwen-32B Qwen2.5-72B
@@ B.2 Model Scale
Methods Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B
@@
Methods Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B

### v4

@@ Abstract
The integration of external knowledge through Retrieval-Augmented Generation (RAG) has become foundational in enhancing large language models (LLMs) for knowledge-intensive tasks. However, existing RAG paradigms often overlook the cognitive step of applying knowledge, leaving a gap between retrieved facts and task-specific reasoning. In this work, we introduce RAG+, a principled and modular extension that explicitly incorporates application-aware reasoning into the RAG pipeline. RAG+ constructs a dual corpus consisting of knowledge and aligned application examples, created either manually or automatically, and retrieves both jointly during inference. This design enables LLMs not only to access relevant information but also to apply it within structured, goal-oriented reasoning processes. Experiments across mathematical, legal, and medical domains, conducted on multiple models, demonstrate that RAG+ consistently outperforms standard RAG variants, achieving average improvements of 3–5%, and peak gains up to 13.5% in complex scenarios. By bridging retrieval with actionable application, RAG+ advances a more cognitively grounded framework for knowledge integration, representing a step toward more interpretable and capable LLMs.
footnotetext: Work done during an internship at Huawei Technologies Ltd.footnotetext: *Corresponding Author
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a broad range of natural language processing tasks (Rong et al., 2025; Tang et al., 2025). To further improve their capabilities, Retrieval-Augmented Generation (RAG) has emerged as a widely adopted framework. By equipping LLMs with access to external knowledge sources, RAG enables the dynamic retrieval of up-to-date information at inference time, leading to significant gains in knowledge-intensive scenarios (Li et al., 2025; Mostafa et al., 2025).
@@
While effective in factual recall and open-domain question answering, RAG frequently underperforms on domain-specific reasoning tasks (Lin et al., 2025; Hayashi et al., 2025; Ammann et al., 2025), where solving complex problems requires not only relevant information but also reasoning about how to use it to arrive at a solution. Figure 1 illustrates a representative example.
RA-DIT Lin et al. (2023) fine-tunes both the retriever and generator in a dual-instruction manner, aligning retrieval more closely with what the model needs to generate accurate responses. Other recent RAG extensions address the gap between retrieval and application by decomposing reasoning tasks into smaller steps and retrieving relevant knowledge for each substep (Singh et al., 2025; Xiong et al., 2025; Zihao et al., 2024). However, they provide limited guidance on how to apply retrieved knowledge. This shortcoming hampers performance in procedural reasoning tasks that require understanding both the solution process and the associated underlying concepts.
@@
Motivated by these insights, we propose RAG+, a simple yet effective extension to the RAG framework that enhances reasoning by bridging retrieval and generation with an application-aware stage. Instead of retrieving only relevant knowledge, RAG+ additionally retrieves examples that demonstrate how the knowledge is applied in practice-such as structured reasoning chains and stepwise solutions-to ground the model’s output in task-relevant usage and to improve reasoning accuracy.
RAG+ builds a dual corpus: one stores domain knowledge, and the other contains automatically generated application instances aligned to each fact. At inference time, the system retrieves relevant knowledge based on the input query and then fetches aligned application examples to provide practical context. This design encourages the model not only to recall factual content but also to produce outputs that follow grounded reasoning patterns based on prior usage. RAG+ is modular and retrieval-agnostic; it can be integrated into any existing RAG pipeline without changing the model architecture or requiring additional fine-tuning.
@@
  • We identify a key limitation of existing RAG systems: the absence of an application-aware step, which impairs reasoning in complex tasks.
@@ 2 Related Work
For retrieval quality, Rewrite-Retrieve-Read (Ma et al., 2023) rewrites queries to better align with document style. $R^{2}AG$ (Ye et al., 2024) uses a retrieval-aware encoder to highlight key signals, while Query2Doc (Wang et al., 2023a) expands queries into pseudo-documents to clarify intent. In terms of corpus construction, GraphRAG (Edge et al., 2024) integrates knowledge graphs to support entity-centric reasoning. To enhance retrieval precision, Reranking (Wang et al., 2024a) and filtering techniques (Wang et al., 2023b; Pickett et al., 2025) further refine retrieved content.
@@ 3.1 Construction Stage
Application Matching: In domains where real-world cases naturally exemplify the use of specific knowledge, each knowledge item is paired with one or more application instances drawn from authentic scenarios, serving as grounded demonstrations that contextualize and concretize the corresponding knowledge. We refer to these instances as application examples. To establish these pairings, we first assign both problems and knowledge items to broad semantic categories using powerful LLMs with temperature sampling and self-consistency voting [^1] The voting mechanism is not essential. To reduce computational cost, a well-designed prompt can serve as an alternative, allowing the alignment of knowledge–application in a single pass.. Specifically, we first classify knowledge items and application instances separately. Then, within each category, we use LLMs to match each knowledge item with its relevant application examples. The resulting pairings are manually refined to ensure accuracy. Next, within each category, relevance selection is conducted by prompting the model to identify the most pertinent knowledge entries for each problem.
@@ 3.2 Inference Stage
During inference, given a test query, RAG+ first retrieves relevant knowledge items from the knowledge corpus based on semantic and structural similarity, using any retrieval method (e.g., dense retrieval, reranking), as shown in Figure 2(b). For each retrieved knowledge item $k$, its corresponding application example $a$, pre-aligned during the construction stage, is retrieved from the application corpus. The pair $(k,a)$ is then incorporated into a predefined prompt template that guides the model with both factual information and procedural cues. This prompt is subsequently fed into the language model for final answer generation.
@@
Method GLM4-9B Qwen2.5-7B DS-Qwen-7B Qwen2.5-14B DS-Qwen-32B Qwen2.5-72B LLaMA3.3-70B
@@ 4.3 Models
Method LLaMA3.1-8B* DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen3-32B Qwen2.5-72B LLaMA3.3-70B
@@ 5 Results
Method LLaMA3.1-8B Qwen2.5-7B DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen2.5-72B LLaMA3.3-70B
@@ 5.4 Impact of Reranking Model
Method Qwen2.5-7B Qwen2.5-14B
@@
Method Qwen2.5-7B Qwen2.5-14B
@@ 5.5 Effect of Application Examples Only
We conducted experiments to assess the contribution of application examples without their associated knowledge on the legal and math domains. Results are shown in Table 6 and Table 7. Incorporating only application examples leads to performance improvements over the base models in most cases. This indicates that application-level guidance alone provides a meaningful benefit. However, this setting still underperforms compared to the full configuration where both knowledge and application examples are included, referred to as RAG+ and AFRAG+.
@@
Method Qwen2.5-7B Qwen2.5-14B DS-Qwen-32B
@@
Table 7: Accuracy (%) on math tasks across different RAG methods. “–” indicates application-only retrieval; “+” indicates the inclusion of retrieved knowledge and application.
Method Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B
Baseline 58.84 66.98 69.07
RAG- 58.19 69.53 69.07
RAG 64.65 73.49 70.47
RAG+ 65.58 74.67 73.72
AFRAG- 59.81 66.98 74.19
AFRAG 63.51 66.98 71.86
AFRAG+ 64.42 70.00 76.05
5.6 Effect of Open-Ended Experiments
Table 8: Accuracy (%) of open-ended sentencing prediction results.
Method Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B
Baseline 49.00 75.50 85.00
RAG 61.50 75.50 89.50
RAG+ 66.50 76.50 90.50
AFRAG 63.00 71.00 83.50
AFRAG+ 66.50 79.50 86.00
We additionally conducted open-ended experiments on the legal sentencing prediction dataset. In this setting, the model is asked to generate a free-form sentence length prediction (e.g., “6–12 months”), and the output is considered correct if it falls within the gold interval. The results are summarized in Table 8. RAG+ consistently improves performance across all model sizes, demonstrating robustness even under open-ended evaluation. This further supports the central hypothesis that combining knowledge with aligned application examples enhances reasoning and task performance.
@@ 6 Discussion
To enable application-aware retrieval, we propose a lightweight method for constructing application-specific corpora from a knowledge base, typically generating one to two examples per item. The corpus grows linearly with the number of items and introduces no retrieval overhead, as each item is directly paired with an application example. In experiments, the initial corpus sizes were 223 KB (math), 528 KB (legal), and 99,382 KB (medicine). Generating the medicine corpus with Qwen2.5-72B on eight 64 GB NPUs took around six hours, which is acceptable given this scale. Final sizes reached 612 KB, 868 KB, and 105,558 KB, respectively, closely matching the sizes of the underlying knowledge corpora. The method is scalable and supports efficient incremental updates.
@@ 7 Conclusion
Acknowledgement
This work was supported by National Natural Science Foundation of China (62232014, 62272377).
@@ Appendix A Experimental Details
For models generating long-form outputs (i.e., QwQ-32B, DS-Qwen-7B, and DS-Qwen-32B), we use stochastic decoding (temperature = 1, top_p = 1) to promote output diversity. For all other models, we apply a deterministic decoding strategy with a temperature of 0 and top_p of 1. Each experiment is conducted three times, and we report the average performance across these runs.
@@ A.3.1 Math QA in Mathematic Domain
We first collected a range of publicly available online resources, supplemented with original questions and corresponding answers. The dataset includes both standard QA items and multiple-choice questions with varying numbers of answer options. Only the correct answers were retained. For compound questions that inquire about multiple values (e.g., "What is the value of $y$ when $x=0$ or x=1x=1?"), we decomposed them into separate single-answer questions, such as "What is the value of $y$ when x=0x=0" and "What is the value of $y$ when x=1x=1?".
Next, we transformed all QA pairs into a multiple-choice format by prompting powerful LLMs to generate three additional plausible but incorrect answer options. This step expands the test set while simultaneously constraining the model’s generative space.
@@ B.1 Complete Results
Method LLaMA3.1-8B* DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen3-32B Qwen2.5-72B LLaMA3.3-70B
@@
Method LLaMA3.1-8B Qwen2.5-7B DS-Qwen-7B DS-Qwen-32B QwQ-32B Qwen2.5-72B LLaMA3.3-70B
@@
Method GLM4-9B Qwen2.5-7B DS-Qwen-7B Qwen2.5-14B DS-Qwen-32B Qwen2.5-72B LLaMA3.3-70B
@@
Method Qwen2.5-14B Qwen3-32B DS-Qwen-32B Qwen2.5-72B
@@ B.2 Model Scale
Method Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B
@@
Method Qwen2.5-7B Qwen2.5-14B Qwen2.5-72B